# Large Language Models as Proxies for Theories of Human Linguistic Cognition

### v1

@@ 3 Large Language Models and alignment with the stimulus
Consider then a theory of HLC - $H_{3}$ above if one so wishes, or something else -that is neutral with respect to some linguistic property $PROP$ for which generative theories include a strong linguistic bias. A potential argument against this relatively linguistically-neutral theory and in favor of a generative theory might take the form of an APS: if humans show knowledge of $PROP$ after exposure to a corpus that seems too impoverished to support the learning of $PROP$ by a linguistically-neutral learner, this will support the strongly biased generative theory. The problem is that it is generally very hard to determine what a given learner might acquire from a given corpus, even for a theory such as $H_{3}$ for which both the representations and the learning mechanism are explicitly provided.
LLMs might be able to help. We can train them on a suitable corpus and then probe their knowledge. Note that this is not entirely straightforward, for various reasons discussed in Lan et al. 2024b. Most immediately, while we might be able to directly inspect a grammar within a theory of HLC such as $H_{3}$ and ask whether it has knowledge of $PROP$, we cannot do the same with current LLMs, whose inner workings are extremely complex and remain mostly opaque, as discussed in Section 2.1. Instead of inspection, we might turn to behavioral measures such as acceptability judgments in minimal pairs where one member satisfies $PROP$ and is acceptable and a minimal variant violates it and is unacceptable, a methodology that has been used in Marvin and Linzen 2018, Hu et al. 2020 and Wilcox et al. 2019, among others. But while the theory of HLC under consideration might offer such judgments, current LLMs provide likelihood assessments, which are an entirely different thing, as mentioned above. However, if we focus on areas where acceptability and likelihood are reasonably well aligned, we might be justified in looking at whether the LLM assigns a higher probability to the $PROP$-satisfying member of the pair than to the $PROP$-violating one. If this probabilistic preference is sufficiently strong, we might consider the fact that the LLM has acquired this preference as reason to think that the learner under our target theory of HLC would acquire the actual $PROP$. This might be so even if the probabilistic preference of the LLM reflects a flawed approximation of $PROP$ rather than knowledge of $PROP$ in any meaningful sense.
@@ 3.1 Methods
bert-based-uncased $\approx$ 3.5 billion tokens 320 years
GPT2 $\approx$ 8 billion tokens 730 years
llama3.2-3b $\approx$ 9 trillion tokens 821,250 years
Table 2: Training data size of the eight language models considered here, and the human linguistic experience equivalent to these data sizes, following Hart and Risley.
Let us illustrate this kind of reasoning with concrete examples. To do so, we use an empirical setup that consists of eight models of different architectures and training schemes (see Section A). [^4] All experimental material can be found in https://osf.io/5zh6q/. The models were trained on datasets that roughly correspond to different timeframes of linguistic experience, ranging from 10 months of human experience (CHILDES) through eight years of linguistic experience (English Wikipedia subset) to 320,730 or even 821,250 years of linguistic experience (BERT base uncased, GPT-2 and Llama 3.2-3b respectively). [^5] More concretely, we use two models from Yedetore et al., an LSTM and a Transformer, that were trained on the CHILDES corpus of child-directed speech; a Transformer we trained on a subset of English Wikipedia, for which we used one of the large Transformer architectures used in Yedetore et al.; the base-strict and base-strict-small versions of the BabyLM LLMs from Warstadt et al. (2023a), trained on the 100M and 10M BabyLM datasets correspondingly; the BERT base uncased version of BERT (Devlin et al. 2018a), trained on the BooksCorpus and English Wikipedia datasets; OpenAI’s GPT-2 (Radford et al. 2019a); Meta’s Llama 3.2-3b LLAMA model (released September 2024). See Section A for further technical detail. The models were probed on their knowledge of three linguistic phenomena using the success criterion introduced above (see Table 3). The linguistic test cases presented in the sections below show that all models, despite being trained on multitudes of data, perform around chance level and worse on the lenient success criterion. This result is not optimistic for LLMs as a proxy for an adequate theory of HLC.
@@
Across-the-board movement (ATB) Which boy did you say that Kim hated and Mary loved yesterday?
* Which boy did you say that Kim hated and Mary loved Ann yesterday?
@@
That-trace effects (TTE) Who did you say that Sue loves?
@@
* Who did you say that loves Sue?
Table 3: Examples of linguistic phenomena with grammatical and ungrammatical sentences. Model probabilities are compared on the underscored critical regions.
@@ 3.2 Test Case: Across-the-board movement and parasitic gaps
The common analysis of wh-movement in generative theories involves a constituent that is attached in one position and is then re-attached higher up in the structure. In (3.2), the constituent [which book] is attached within the embedded clause as a sister to Mary and is then re-attached at or near the root of the matrix clause, which among other things leads to it being pronounced in the beginning of the sentence.
Some work within generative linguistics (Pearl and Sprouse 2013, Phillips 2011) suggests that a relatively linguistically-neutral theory cannot account for the full knowledge of wh-movement given the kind of data that children are exposed to, an instance of the APS. This is not a consensus view, however, and some recent work such as Wilcox et al. suggests that LLMs undermine this APS and that these models show that a linguistically neutral theory would, in fact, acquire a full knowledge of wh-movement from a developmentally-realistic corpus. Following Lan et al. let us show how LLMs (used as proxies for an undisclosed linguistically-neutral theory of HLC) might bear on the APS under consideration and why they currently do nothing to undermine it.
Before proceeding, let us reiterate that it is hard to tell whether LLMs are good proxies for the relatively linguistically-neutral theory of HLC that is envisioned in works such as Wilcox et al. (2023) in the absence of at least some specifics about that theory. But if they are, and if we focus on cases where correctness and likelihood are reasonably well-aligned, we could reason that if the LLM overwhelmingly assigns much higher probability to the correct member of each minimal pair, this indicates that there was sufficient information in the training data for the target linguistically-neutral theory to have acquired the relevant knowledge. And if the LLM does not perform as well, this suggests that the target theory would also struggle. The former possibility weakens the APS, and the latter strengthens it.
@@
We test four models from Table 2 on their knowledge of across-the-board movement and parasitic gaps using the minimal pair method. We compare sentences that both have an upstream filler - like which book in Example 3.2, or who in Example 3.2 - and differ in whether they are gapped accordingly. [^6] The paradigm sentences in Lan et al. (2024b) were generated by template for ATB and PG. Here we unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names), and by adding more lexical choices. From each template, represented by a context-free grammar, we sampled 10,000 sentence pairs and ran them through each model to get the relevant probability values (see Appendix B). As is shown in Figure 1, all prefer the ungrammatical over the grammatical members of the minimal pairs in the vast majority of cases. This failure suggests that the linguistically-neutral theory for which the LLMs are taken to be useful proxies would not acquire wh-movement from a developmentally-realistic corpus.
Figure 1: Model accuracy on ATB and PG datasets averaged over five experiment seeds. Accuracy is measured as the ratio of cases where the model assigns a higher probability to the grammatical sentence continuation.
3.3 Test Case: that-trace effects
Another property of wh-movement in English is that when the moved wh-phrase is a subject, the clause from which it moves cannot have an overt complementizer such as ‘that’:
@@ 3.3 Test Case: that-trace effects
This asymmetry, known as that-trace effects (TTE), has been thoroughly discussed in theoretical linguistics (see Perlmutter 1968, Chomsky and Lasnik 1977, and Phillips 2013a, among others). It has been shown that extraction of any kind from clauses with that is vanishingly rare in adult-directed corpora, and even rarer in child-directed speech Phillips (2013a) [^7] Out of 11,308 analyzed utterances from child-directed speech, there were only two instances of non-subject extraction with that and zero instances of subject extraction with that, contrasting with 159 instances of object extraction without that, and 13 instances of subject extraction without that.. This distribution is claimed (Chomsky and Lasnik 1977, Phillips 2013b) to be insufficient to warrant acquisition of the paradigm above. We set aside the important theoretical discussion of what bias present in the architecture or representations of $H_{1}$ underlies the TTE according to generative linguistics. [^8] See Sobin (1987), Lohndal (2009), Bošković (2016) and Pesetsky (2017), among others, for discussion. Instead, we focus on its status as an APS, which warrants evaluating the performance of the linguistically-neutral LLM contender on it.
Similarly to the previous case, we test eight LLM models (see Table 2). The four-way asymmetry warrants two criteria that constitute knowledge of the TTE: first, given an overt complementizer upstream, the model should prefer an overt subject to a gap in the embedded clause. This preference should translate into probabilities so that $P(loves)$ in 3.3 should be smaller than $P(Sue)$ in 3.3. Second, given a gapped sentence downstream ("…       loves Sue"), the model should prefer a prefix without an overt complementizer. That is, $P(loves)$ in 3.3 is expected to be smaller than $P(loves)$ in 3.3. This way, we measure the dispreference for an overt complementizer in the presence of a moved subject along both axes - that of the complementizer and that of the gap.
Results for both success criteria are presented in Figure 2. Under both criteria, the performance of the models hovers around chance. While it is unclear how to interpret these results, we see nothing in the performance of the models that suggests that the target hypothesis will account for the learning of that-trace effects from similar training corpora.
Figure 2: Model accuracy values for the $P(+that,+trace)<P(+that,-trace)$ criterion (top) and $P(+that,+trace)<P(-that,+trace)$ criterion (bottom), over samples of 10,000 sentence pairs from the TTE test set. Results are averaged over five seeds. The dark plot represents model training sizes.
4 Large Language Models and cross-linguistic evidence
Consider again the comparison of a strongly linguistically-biased generative theory and a more linguistically-neutral alternative. Suppose that the strong linguistic biases of the generative theory can be shown to help derive a cross-linguistic pattern, perhaps by making it easier to represent and learn the languages that adhere to the pattern than those that do not. In the linguistically-neutral theory it can be much harder to see directly whether the attested patterns are predicted to be easier to learn than the unattested ones. One might try to evaluate this matter empirically, by simulating learning of languages of the attested and the unattested kind given the linguistically-neutral theory, but this can be hard to do. Consequently, it can be hard to see whether the theory can explain the cross-linguistic pattern under consideration and whether it can compete with the generative theory. Again, however, if an LLM is a good proxy for the theory, the problem is avoided to some extent. One can examine the ease with which the LLM learns to approximate both languages that follow the pattern and languages that violate it. If the LLM approximates the former more easily than the latter, this can suggest that the target theory might be able to account for the cross-linguistic pattern (subject to various further assumptions about how relative ease of learning derives the typology, and to an adequate definition of how ease-of-learning via LLM is to be measured). If, on the other hand, some of the pattern-violating languages are easier to learn than some of the pattern-satisfying ones, the linguistically-neutral theory would need further mechanisms to derive the typology.
We should emphasize that the inferences from this argument are particularly weak. On the one hand, even if the linguistically-neutral theory derives a learning asymmetry between pattern-satisfying and pattern-defying languages but the difference is very small, this might not suffice to account for a robust typological pattern. On the other hand, even if the linguistically-neutral theory does not derive a learning asymmetry, it might be able to account for it in terms of other factors, such as communicative pressure. In order to overcome these limitations one could try to situate the learning component within a model of cultural evolution, which allows for the amplification of small asymmetries over generations and also factors in communicative pressures and other considerations. We believe that such a model is the proper context for the use ease of learning by LLMs to evaluate typological asymmetries, but following Kallini et al. (2024), whose setup we use here, we leave such a model for future work and only assess currently available models. [^9] Kallini et al. (2024) contextualize their work differently. We will not attempt to determine whether the present work is in line with their goals.
4.1 Ease of learning by LLM as a theory of HLC?
Kallini et al. (2024) compare GPT-2 when trained on an English corpus and when trained on corpora that are derived from the original corpus through various perturbations, each corresponding to an unattested language. They find that the original English corpus was easier for GPT-2 than the derived corpora of unattested languages when ease-of-learning is measured through test set perplexities through time (see Figure 3). In doing so, Kallini et al. (2024) make use of LLMs as proxies in the context of explaining the cross-linguistic skew that rules out the unattested perturbations of English, even though this theory is never made explicit in the paper. If a model like GPT-2 is indeed a good proxy for this contender, then we may say that what Kallini et al. (2024) have shown to hold for English should also hold cross-linguistically, as presented in the introduction: all unattested languages should be harder to learn than attested ones.
In the following subsections we consider a few languages for which the evaluation is simple given current resources and test them with the perturbations used by Kallini et al. (2024) and one additional perturbation, listed in Table 4. For each perturbation, we create perturbed versions and baseline versions for each of the four languages - English, Italian, Russian and Hebrew, based on datasets from Gulordava et al. (2018). We then train a transformer LLM based on the architecture and training scheme from Yedetore et al. (2023), and consider the validation set perplexities as training progresses. [^10] The model was trained for 48 hours and validation perplexity was computed every 200 training batches. The 48 hours of training amounted to roughly 1.5 epochs per dataset, with batch size 10, such that each epoch amounted to an average of 181,552 batches, see Appendix B for further details. We find we find several different languages of an unattested kind that are easier for the LLM to approximate than a corresponding attested language, as measured by the perplexity metric used by Kallini et al. (2024). From the perspective of the Proxy View, this means that the target, relatively linguistically-neutral theory fails to receive support from the LLM: if the target hypothesis can explain the typological pattern, this is not detected by the LLM experiment.
Perturbation Attested / Perturbed
partial-reverse Colorless0 green1 <rev>2 ideas3 sleep4 furiously5.
Colorless0 green1 <rev>2 furiously5 sleep4 ideas3.
full-reverse Colorless0 green1 <rev>2 ideas3 sleep4 furiously5.
Furiously5 sleep4 ideas3 <rev>2 green1 colorless0.
switch-indices Colorless0 green1 ideas2 sleep3 furiously4.
Ideas2 green1 colorless0 sleep3 furiously4.
token-hop They0 were1 sleeping2 v3 next4 to5 the6 colorless7 green8 ideas9.
They0 were1 sleeping2 next3 to4 the5 v6 colorless7 green8 ideas9.
Table 4: Perturbation test cases. Ease-of-learning is evaluated for the attested and perturbed versions of English, Italian, Hebrew and Russian.
4.2 Test case: partial-reverse
We demonstrate the ease-of-learning method in our first test case - a cross-linguistic reproduction of the partial-reverse perturbation from Mitchell and Bowers 2020 and Kallini et al. 2024. Each sentence is reversed starting from a randomly chosen index in the input sentence, in which a special marker token <rev> is inserted:
@@ 4.2 Test case: partial-reverse
The target, relatively linguistically-neutral theory in this case might be a version of a generative theory in which there is a parameter for reverse. Any grammar that is statable within the theory now comes in two variants. If the parameter is set to 0, the grammar works as in the original generative theory. But if the parameter is set to 1, every derivation ends with reversing the output starting from a random point. This may or may not be the target theory that Kallini et al. have in mind; other target theories are of course imaginable, and Kallini et al. do not comment explicitly about this. We note that, like Kallini et al.’s other perturbations, reversal seems quite far away from the typological questions that generative linguistics have usually concerned themselves with. Here we stay with Kallini et al.’s perturbations and setup simply in order to illustrate the relevant methodological considerations, but a proper exploration of the Proxy View would need to look at those cross-linguistic patterns that have informed linguistic research.
@@
We find that validation perplexities follow an attested-unattested divide: the partial-reverse perturbed versions of English, Italian and Russian are harder to learn than their attested counterparts (see Fig. 3). [^11] The partial-reverse languages are compared to a modified baseline of English in which the marker <rev> is inserted randomly without reversing, to control for the effect for additional textual material in the sentence on perplexities. Since we use the same seed for both perturbations, the markers are inserted in the same indices (see example 4.2).
This, as discussed above, could be part of an argument that linguistic biases are not needed to rule out partial-reverse languages, but crucially only if there were a serious contender for the best explanation that allows for such languages in the first place, and could be compared theoretically with the restrictive generative account. However, other perturbations paint a different picture under the cross-lingustic view.
4.3 Test case: full-reverse
We now consider the perturbation $\Pi=$ full-reverse from Kallini et al. 2024, where each sentence is reversed in its entirety and a special marker token is randomly inserted:

### v2

@@ 2 Large language models and alignment with the stimulus
Consider then an $H_{3}$ that is neutral with respect to some linguistic property $PROP$ for which generative theories such as $H_{1}$ include a strong linguistic bias. A potential argument against $H_{3}$ and in favour of $H_{1}$ might take the form of an APS: if humans show knowledge of $PROP$ after exposure to a corpus that seems too impoverished to support the learning of $PROP$ by a linguistically-neutral learner, this will appear to support the strongly biased $H_{1}$. The problem is that it is generally very hard to determine what a given learner might acquire from a given corpus, even for a theory such as $H_{3}$ for which both the representations and the learning mechanism are explicitly provided.
LLMs might be able to help. We can train them on a suitable corpus and then probe their knowledge. Note that this is not entirely straightforward, for various reasons discussed in Lan et al. 2024b. Most immediately, while we might be able to directly inspect a grammar within a theory of HLC such as $H_{3}$ and ask whether it has knowledge of $PROP$, we cannot do the same with current LLMs, whose inner workings are extremely complex and remain mostly opaque, as discussed in Section 1.1. Instead of inspection, we might turn to behavioural measures such as acceptability judgments in minimal pairs (see Section 1.1 above), where one member satisfies $PROP$ and is acceptable and a minimal variant violates it and is unacceptable. An immediate concern with this methodology is that, while the theory of HLC under consideration might offer such judgments, current LLMs provide only likelihood assessments, through their distribution over the next token given a prefix. Likelihood assessments are very different from acceptability judgments, as mentioned above. However, if we focus on areas where acceptability and likelihood are reasonably well aligned we might be justified in looking at whether the LLM assigns a higher probability to the $PROP$-satisfying member of the pair than to the $PROP$-violating one (a methodology that has been used in Marvin and Linzen, 2018, Hu et al., 2020 and Wilcox et al., 2019, among many other works; see Section 1.1). If this probabilistic preference is sufficiently strong, we might consider the fact that the LLM has acquired this preference as reason to think that the learner under our target theory of HLC $H_{3}$ would acquire the actual $PROP$. This might be so even if the probabilistic preference of the LLM reflects a flawed approximation of $PROP$ rather than knowledge of $PROP$ in any meaningful sense.
@@ 2.1 Methods
BERT Base Uncased $\approx$ 3.5 billion tokens 320 years
GPT-2 $\approx$ 8 billion tokens 730 years
LLaMA-3.2-3b $\approx$ 9 trillion tokens 821,250 years
Таблица 2: Training data size of the eight language models considered here, and the human linguistic experience equivalent to these data sizes.
Let us illustrate this kind of reasoning with concrete examples. To do so, we use an empirical setup that consists of eight models of different architectures and training schemes, ranging from the rough equivalent of several months of human linguistic input to hundreds of thousands of years, with multiple models in between, as summarized in Table 2.555We use two models from Yedetore et al. 2023, an LSTM and a Transformer, that were trained on the CHILDES corpus of child-directed speech; a Transformer we trained on a subset of English Wikipedia, for which we used one of the large Transformer architectures used in Yedetore et al. 2023; the base-strict and base-strict-small versions of the BabyLM LLMs from Warstadt et al. (2023), trained on the 100M and 10M BabyLM datasets respectively; the BERT base uncased version of BERT (Devlin et al., 2018), trained on the BooksCorpus and English Wikipedia datasets; OpenAI’s GPT-2 (Radford et al., 2019) and Meta’s LLaMA 3.2-3b LLaMA model (released September 2024, Meta AI, 2024). See Appendix A for further technical detail. All experimental material can be found at https://osf.io/5zh6q/. Human equivalents are based on the assumption by Hart and Risley (1995) that children are exposed to around 30,000 words per day, while roughly equating words with model tokens.
The models were tested on three linguistic phenomena using the very lenient success criterion discussed above (see Table 3). The models that are most directly relevant to assessing $H_{3}$ with respect to any given APS are the smaller ones, which are trained on inputs that roughly match the number of words that a child might be exposed to. If these models prefer the grammatical option over the ungrammatical one in the minimal pairs of the test set for a given phenomenon, this can support the idea that $H_{3}$ would actually acquire the relevant piece of linguistic knowledge from a developmentally-realistic corpus. [^6] Recall that we do not ask whether the LLMs actually acquire the linguistic knowledge under consideration, nor do we expect them to. Conversely, lack of such a preference can weaken the plausibility that $H_{3}$ would succeed in acquiring the relevant knowledge from a similar corpus. [^7] As should be clear, even if there is evidence that $H_{3}$ could acquire the relevant knowledge from a developmentally-realistic corpus there might still be good reasons to adopt an $H_{1}$ that is strongly biased in favour of this knowledge. Among other things, $H_{1}$ might provide a better explanation than $H_{3}$ for the cross-linguistic pattern. We believe that this is the case for the phenomena discussed in the present section and refer the reader to the literature cited below.
Two of the models (BERT and GPT-2) were trained on the equivalent of hundreds of years of linguistic input. This is clearly beyond what a human child has been exposed to by the time they show evidence of the relevant linguistic knowledge. We included these models to allow proponents of the Proxy View to use LLMs as broadly adequate proxies for $H_{3}$ that just happen to be extremely data inefficient.
We included the much larger LLaMA mainly to see whether current architectures are sensible proxies for $H_{3}$ for the phenomena at hand. If they cannot show even a slight preference for the correct element in minimal pairs after training on the equivalent of hundreds of thousands of years of linguistic input, then perhaps they are incapable of approximating the pattern under consideration, which would suggest that they are not very good proxies for $H_{3}$ with respect to this pattern.
@@
Across-the-board movement (ATB) Which boy did you say that Kim hated and that Mary loved yesterday?
* Which boy did you say that Kim hated and that Mary loved Ann yesterday?
@@
That-trace effects (TTE) * Who did you say that loves Sue?
Who did you say that Sue loves?
@@
Таблица 3: Examples of linguistic phenomena with grammatical and ungrammatical sentences. Model probabilities are compared on the underscored critical regions.
@@ 2.2 Test case: across-the-board movement and parasitic gaps
Above we mentioned displacement phenomena, illustrating with the position of the auxiliary in yes-no questions in English. Displacement phenomena have featured in recent discussions of the role of LLMs in evaluating APSs, both for the placement of the auxiliary (see, e.g., Yedetore et al., 2023) and for wh-movement (see Chowdhury and Zamparelli, 2018, Bhattacharya and van Schijndel, 2020, Chaves, 2020, Warstadt et al., 2020, Ozaki et al., 2022, Wilcox et al., 2023, and Lan et al., 2024b). The latter will concern us in the current two subsections. Here is a simple illustration:
@@
[Which book] did you say that Mary read    last week?
The common analysis of wh-movement in generative theories involves a constituent that is attached in one position and is then re-attached higher up in the structure. In (2.2), the constituent [which book] is attached within the embedded clause as a sister to read and is then re-attached at or near the root of the matrix clause, which among other things leads to it being pronounced in the beginning of the sentence.
Some work within generative linguistics (Pearl and Sprouse, 2013, Phillips, 2013) suggests that a relatively linguistically-neutral theory cannot account for the full knowledge of wh-movement given the kind of data that children are exposed to, an instance of the APS. This is not a consensus view, however, and some recent work such as Wilcox et al. 2023 suggests that LLMs undermine this APS and that these models show that a linguistically-neutral theory would, in fact, acquire a full knowledge of wh-movement from a developmentally-realistic corpus. Following Lan et al. let us show how LLMs (used as proxies for an undisclosed linguistically-neutral theory of HLC) might bear on the APS under consideration and why they currently do nothing to undermine it.
Before proceeding, let us reiterate that it is hard to tell whether LLMs are good proxies for the relatively linguistically-neutral $H_{3}$ that seems to be envisioned in works such as Wilcox et al. 2023 in the absence of at least some specifics about that theory. But if they are, and if we focus on cases where correctness and likelihood are reasonably well-aligned, we could reason that if the LLM overwhelmingly assigns much higher probability to the correct member of each minimal pair, this indicates that there was sufficient information in the training data for the target linguistically-neutral theory $H_{3}$ to acquire the relevant knowledge. And if the LLM does not perform as well, this suggests that the target theory $H_{3}$ might also struggle. The former possibility weakens the APS, and the latter strengthens it.
Following Lan et al. (2024b) we focus on a family of wh-movement configurations where there is complex interaction between two gaps, in a way that seemingly violates island constraints on syntactic movement. [^8] See Ross 1967, Williams 1977, 1990, Engdahl 1983, Haïk 1985, Munn 1992, Postal 1993, Fox 2000, Nissenbaum 2000, Hornstein and Nunes 2002, and Sportiche 2024, among many others, for many important observations about the relevant phenomena that we will not be able to discuss in the present paper. Our simplified treatment here suffices to illustrate the methodological point about evaluating the Proxy View, and since our conclusions below are negative even for the simplified picture, we see no reason to think that current models will successfully approximate the complexities of these phenomena discussed in the literature. Any attempt to make a positive claim about these phenomena would of course need to address these complexities. The first such phenomenon is across-the-board (ATB) movement in coordinate structures. While extraction from a single coordinate is ungrammatical, as in 2.2 and 2.2, extraction can become grammatical when it takes place from both coordinates, as in 2.2:
@@
The second nuance of wh-movement we test is when extraction from an island is made possible by the presence of another gap downstream, a phenomenon known as a parasitic gap (PG):
@@
We test the eight models from Table 2 on their approximation of across-the-board movement and parasitic gaps using the minimal pair method. We compare sentences that both have an upstream filler – like which book in Example 2.2, or who in Example 2.2 – and differ in whether they are gapped accordingly. [^9] The paradigm sentences in Lan et al. 2024b were generated by separate templates for ATB and PG. Here we unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names), and by adding more lexical choices. From each template, represented by a context-free grammar, we sampled 10,000 sentence pairs and ran them through each model to get the relevant probability values (see Appendix B). As shown in Figure 1, all models other than LLaMA prefer the ungrammatical over the grammatical members of the minimal pairs in the vast majority of cases. This failure suggests that the linguistically-neutral $H_{3}$ for which the LLMs are taken to be proxies would not acquire wh-movement from a developmentally-realistic corpus.
Рис. 1: Model accuracy on ATB and PG datasets averaged over five experiment seeds. Accuracy is measured as the proportion of cases where the model assigns a higher probability to the grammatical sentence continuation. The dark plot represents model training sizes (on a logarithmic scale, right).
2.3 Test case: that-trace effects
Another property of wh-movement in English is that when the moved wh-phrase is a subject, the clause from which it moves cannot have an overt complementizer such as that:
@@ 2.3 Test case: that-trace effects
This phenomenon, known as that-trace effects (TTE), has been the topic of much discussion in the linguistic literature. [^10] See Perlmutter 1968, Bresnan 1972, Chomsky and Lasnik 1977, and much subsequent work. See Pesetsky 2017 for a more recent overview and discussion. As with ATB movement and PG, the present discussion is a simplification, and we refer the reader to the literature for observations and analysis that would be important for any attempt to make positive claims about LLM performance and its potential support for $H_{3}$. Building on corpus data from Pearl and Sprouse 2013, Phillips (2013) points out that extraction of any kind from clauses with that is vanishingly rare in adult-directed corpora and even rarer in child-directed speech: out of 11,308 analyzed utterances from child-directed speech, there were only two instances of non-subject extraction with that and zero instances of subject extraction with that, contrasting with 159 instances of object extraction without that, and 13 instances of subject extraction without that. As Phillips (2013) notes, these distributional observations strongly suggest an APS for the knowledge behind TTE. Still, one might try to defend an $H_{3}$ that is not biased in favour of TTE and in which knowledge of TTE is acquired from the input data, perhaps pieced together from several indirect sources of evidence. As with the ATB and PG cases above, we use LLMs as proxies for such a linguistically-neutral $H_{3}$: if these models adequately approximate TTE, this can be taken to suggest that $H_{3}$ might succeed in learning TTE, thus weakening the current APS.
@@
Results for the $P(\text{+trace}|\text{+that})<P(\text{-trace}|\text{+that})$ criterion are presented in Figure 2. All models except LLaMA overwhelmingly prefer the ungrammatical element to the grammatical one in the minimal pairs of the test. This is very similar to what we saw for ATB and PG above. It is thus hard for us to see how these results can be used to support the ability of $H_{3}$ to acquire knowledge of TTE from a developmentally-realistic corpus. [^11] As a reviewer notes, our results are in tension with the apparent ability of more recent versions of ChatGPT to handle TTE (see Pesetsky, 2024). In the absence of sufficient detail by OpenAI about its training process we cannot pinpoint the source of this tension, but we note that the training corpora for current versions of GPT are many orders of magnitude bigger than those used for our models, with the possible exception of LLaMA. As the performance of LLaMA reported here suggests, training on sufficiently large corpora (the equivalent of hundreds of thousands of years of linguistic exposure) may suffice for above chance performance by current models. Moreover, it is conceivable that the pretraining of these current versions includes academic papers that explicitly discuss TTE. It is also possible that at least some of the apparent awareness of these versions to TTE comes from the human-guided stages that follow pretraining.
Рис. 2: Model accuracy for the $P(\text{+trace}\mid\text{+that})<P(\text{-trace}\mid\text{+that})$ criterion over samples of 10,000 sentence pairs from TTE test set. Results are averaged over five seeds. The dark plot represents model training sizes (on a logarithmic scale, right).
3 Large language models and cross-linguistic evidence
Consider again the comparison of a strongly linguistically-biased generative theory $H_{1}$ and a more linguistically-neutral alternative $H_{3}$. Suppose that the strong linguistic biases of the generative theory $H_{1}$ can be shown to help derive a cross-linguistic pattern, perhaps by making it easier to represent and learn the languages that adhere to the pattern than those that do not. In the relatively linguistically-neutral theory $H_{3}$ it can be much harder to see directly whether the attested patterns are predicted to be easier to learn than the unattested ones. One might try to evaluate this matter empirically, by simulating learning of languages of the attested and the unattested kind given the linguistically-neutral theory, but this can be hard to do. Consequently, it can be hard to see whether the theory can explain the cross-linguistic pattern under consideration and whether it can compete with the generative theory. Again, however, if an LLM is a good proxy for $H_{3}$, the problem is mitigated to some extent. One can examine the ease with which the LLM learns to approximate both languages that follow the pattern under consideration and languages that violate it. If the LLM approximates the former more easily than the latter, this can suggest that the target theory might be able to account for the cross-linguistic pattern (subject to various further assumptions about how learning affects the typology). If, on the other hand, some of the pattern-violating languages are easier to learn than some of the pattern-satisfying ones, the linguistically-neutral theory would need further mechanisms to derive the typology. We outline a preliminary investigation of this kind, building on the setup of Kallini et al. (2024). We discuss limitations of this approach in Section 3.6.
3.1 Ease of learning by LLM and typological asymmetries
Kallini et al. (2024) compare GPT-2 when trained on an English corpus and when trained on corpora that are derived from the original corpus through various perturbations, each perturbed corpus corresponding to an unattested kind of language. Kallini et al. find that the original English corpus was easier for GPT-2 than the derived corpora of languages of unattested kinds on a very particular notion of ease of learning, defined through test set perplexities through time (see Figure 3). In doing so, Kallini et al. can be taken to make use of LLMs as proxies for a linguistically-neutral $H_{3}$ in the context of explaining various cross-linguistic skews. These skews might initially be taken to support an $H_{1}$ that is strongly linguistically biased against languages corresponding to the perturbations under consideration. The performance of GPT-2, however, might suggest that a more linguistically-neutral $H_{3}$ is also compatible with the typology. On this view, languages corresponding to the perturbations might be unattested because of general, non-linguistic biases that are shared by $H_{3}$ and GPT-2.121212Kallini et al. (2024) contextualize their work differently, relating it to the notion of impossible languages (Moro, 2016; Musso et al., 2003). We will not attempt to determine whether the present work is in line with their goals.
In the following subsections we extend the setup of Kallini et al. 2024 by adding to their examination of English a few other languages for which the evaluation is simple given current resources. We use corpora for these languages as the basis for the perturbations used by Kallini et al. and for one additional perturbation, listed in Table 4. For each perturbation, we use a baseline and a perturbed version for each of the four languages – English, Italian, Russian, and Hebrew, based on datasets from Gulordava et al. 2018. We then train a transformer LLM based on the architecture and training scheme from Yedetore et al. (2023) and consider the validation set perplexities as training progresses. [^13] The model was trained for 48 hours and validation perplexity was computed every 200 training batches. The 48 hours of training amounted to roughly 1.5 epochs per dataset, with batch size 10, such that each epoch amounted to an average of 181,552 batches. See Appendix B for further details.
As outlined above, if the setup in Kallini et al. 2024 is intended to support $H_{3}$, then the behaviour of GPT-2 with respect to their perplexity metric should show that $H_{3}$ can be part of a successful account of the typology and that the strong linguistic biases of $H_{1}$ are not necessary. A relatively straightforward argument would be if ease of learning directly matched the typological divide across languages, with languages of the attested kind being uniformly easier to learn than languages of the unattested kind. But ease of learning may well be just one out of several interacting factors that give rise to the typology; other plausible factors include communicative efficiency, sociolinguistic pressures, and historical accident. Kallini et al.’s setup does not incorporate such factors explicitly, but one can perhaps consider them as part of the background. Even when such factors are taken into consideration, however, a setup such as Kallini et al. 2024’s would presumably only be useful if it showed that ease of learning by the relatively linguistically-neutral $H_{3}$ can play some role in an account of the typology. Perhaps, rather minimally, the Kallini et al. approach could succeed within languages, in that the linguistically-unnatural perturbations of each language would be harder to learn than the attested baseline.
As we show, LLM performance does not provide this kind of support to $H_{3}$. Ease of learning in the sense of Kallini et al. (2024) does not match the general typological divide across languages, and even within languages, the perturbed version is sometimes easier to learn than the attested baseline. From the perspective of the Proxy View, this means that the relatively linguistically-neutral $H_{3}$ fails to receive support from the performance of the proxy LLM: if $H_{3}$ can be part of an explanation of the typological pattern, this is not shown by the current results.
Perturbation Attested / Perturbed
partial-reverse (Section 3.2) colourless0 green1 <rev>2 ideas3 sleep4 furiously5.
colourless0 green1 <rev>2 furiously5 sleep4 ideas3.
full-reverse (Section 3.3) colourless0 green1 <rev>2 ideas3 sleep4 furiously5.
Furiously5 sleep4 ideas3 <rev>2 green1 colourless0.
switch-indices (Section 3.4) colourless0 green1 ideas2 sleep3 furiously4.
Ideas2 green1 colourless0 sleep3 furiously4.
token-hop (Section 3.5) They0 were1 sleeping2 v3 next4 to5 the6 colourless7 green8 ideas9.
They0 were1 sleeping2 next3 to4 the5 v6 colourless7 green8 ideas9.
Таблица 4: Perturbation test cases. Ease of learning is evaluated for the attested (top, green) and perturbed (bottom, orange) versions of English, Italian, Hebrew and Russian.
3.2 Test case: partial-reverse
We demonstrate the use of ease of learning in our first two test cases, in which we consider reversed constructions. In our first test case of this sort – a cross-linguistic reproduction of the partial-reverse perturbation from Kallini et al. 2024 – each sentence is reversed starting from a randomly chosen index in the input sentence, in which a special marker token <rev> is inserted: